### v1

# EarthMarker: A Visual Prompt Learning Framework for Region-level and Point-level Remote Sensing Imagery Comprehension

Abstract
Recent advances in visual prompting in the natural image area have allowed users to interact with artificial intelligence (AI) tools through various visual marks such as box, point, and free-form shapes. However, due to the significant difference between the natural and remote sensing (RS) images, existing visual prompting models face challenges in RS scenarios. Moreover, RS MLLMs mainly focus on interpreting image-level RS data and only support interaction with language instruction, restricting flexibility applications in the real world. To address those limitations, a novel visual prompting model named EarthMarker is proposed, which excels in image-level, region-level, and point-level RS imagery interpretation. Specifically, the visual prompts alongside images and text instruction input into the large language model (LLM), adapt models toward specific predictions and tasks. Subsequently, a sharing visual encoding method is introduced to refine multi-scale image features and visual prompt information uniformly. Furthermore, to endow the EarthMarker with versatile multi-granularity visual perception abilities, the cross-domain phased learning strategy is developed, and the disjoint parameters are optimized in a lightweight manner by leveraging both the natural and RS domain-specific knowledge. In addition, to tackle the lack of RS visual prompting data, a dataset named RSVP featuring multi-modal fine-grained visual prompting instruction is constructed. Extensive experiments are conducted to demonstrate the proposed EarthMarker’s competitive performance, representing a significant advance in multi-granularity RS imagery interpretation under the visual prompting learning framework. Our code and dataset are available at https://github.com/wivizhang/EarthMarker.
Index Terms:
Visual prompting, Remote sensing, Multi-modal large language models (MLLMs).I Introduction
Visual prompting refers to the technique of guiding the visual models to focus on the region of interest and improving their finer-grained interaction performance by providing them with visual marks (e.g., boxes, points, masks) or examples[1, 2, 3]. Recently, multi-modal large language models (MLLMs) [4, 5] have experienced remarkable advancements in the remote sensing (RS) domain. However, those MLLMs only support language instruction and fail to understand the images in a visual prompting manner. Considering that the high-resolution RS imagery is characterized by scale variation, across categories and tiny objects, fine-grained reasoning is necessary alongside holistic scene interpretation. This is crucial to perform more detailed analyses to make informed decisions in real-world applications[6]. Nevertheless, most existing MLLMs achieve visual-language alignment using image-text pairs, lacking fine-grained referring understanding abilities, such as region-level and point-level. At present, leveraging the visual prompting method to enhance the complex visual reasoning capabilities of MLLMs in RS remains under-explored.
Notably, prompting engineering [7, 8] has been extensively studied in the natural language processing (NLP) community[9] and subsequently spread to the computer vision area. A key example is the Segment Anything (SAM) [10] model, which utilizes multiple visual prompting marks to realize zero-shot segmentation adapted for various new image distributions. However, SAM lacks semantic information, hindering real-world applications. GPT4RoI [11] and RegionBlip [12] have enabled MLLMs to complete region-level visual understanding tasks by training on region-text pairs. Nevertheless, they only support bounding boxes as visual prompts, which lacks flexibility. Osprey [13] excels in pixel-level visual understanding but relies on pre-attached segmentation models, constraining its application range. Additionally, Ferret [14] and SPHINX-V [15] support free-shape visual prompting marks to achieve pixel-level image comprehension. However, all these models are trained on natural scene data, leading to inferior performance when handling the RS imagery.
In the RS field, there are a few works that limited works were devoted to region-level imagery interpretation. For example, RSVG [16] adopts language prompting to inquire and localize the specific object, but does not involve the visual prompting technique to realize various visual tasks. In addition, inspired by SAM, RSPrompter [17] introduces an automated prompts generation to develop interactive segmentation in RS imagery. Another representative work is EarthGPT [5], which shows the potential of region-level image comprehension by training on visual grounding datasets. However, EarthGPT only supports language interaction without visual prompts, thereby lacking flexibility. These limitations hinder the development of fine-grained spatial understanding and complex reasoning execution in the RS domain. It is clear that the visual prompts learning in the RS domain is still in its infancy.
To bridge this gap, a fine-grained MLLM named EarthMarker is proposed, leveraging visual prompting to extend the capability of MLLMs for region-level and point-level understanding in the RS domain for the first time. Based on the visual prompts learning, as illustrated in Fig. 1, EarthMarker excels at the multi-granularity interpretation of RS imagery across image, region, and point levels. Moreover, EarthMarker can complete a wide range of RS visual reasoning tasks, including scene classification, referring object classification, captioning, inter-relationship analysis, etc. Concretely, the visual prompts, i.e., bounding boxes and points, along with the RS images and the text instructions are provided as input to the LLM. Notably, the visual prompting marks are utilized to isolate specific areas and guide the model to interpret regional content in the entire RS image. Considering that the RS imagery is gathered from an overhead perspective by satellites, associated with large-scale variations and cluttered backgrounds, multi-resolution image input processing is necessary. Subsequently, unlike most existing nature scene visual prompting works, which routinely set different visual encoders and visual prompts encoders. In our method, a sharing visual encoding method is developed. Specifically, the visual prompt is processed to RGB images analogously, which shares the same visual encoder with the inputted image. This strategy is beneficial for consistent feature extraction and understanding the relationship between visual prompts regions and the holistic image, enhancing the performance of the model under visual prompts learning.
In order to enhance the visual prompts-image-text alignment and to equip the EarthMarker with versatile multi-granularity visual comprehension abilities, the cross-domain phased learning strategy is proposed. In the first stage for multi-domain image-text alignment, EarthMarker is trained on the existing nature scene and RS caption data to obtain general image understanding and enhance the modeling of conceptual diversity. Subsequently, the model is further trained on the nature domain referring data to achieve spatial perception in images, beneficial for subsequent developing referring comprehension ability in the RS domain. Lastly, in the RS visual prompting tuning stage, leveraging RS region-text and point-text instruction data, the proposed EarthMarker is equipped with point-level and region-level RS imagery interpretation capability. Notably, the phased training leverages the natural domain generalized knowledge and the RS domain expert knowledge for developing RS visual prompting MLLM. The multi-domain joint training is advantageous for enhancing the deep interpretation of fine-grained RS imagery and improving open-vocabulary reasoning capabilities. In addition, the updatable parameters of the model are disjoint, preventing interference between understanding images at different granularity and the capability to follow the visual prompt instruction.
Another challenge lies in the datasets, e.g., existing visual prompting datasets [14, 13] are restricted to the natural scene, lacking RS semantics. It has become indispensable to construct a visual prompting dataset tailored to the RS domain for developing fine-grained MLLM. To this end, a RS visual prompting dataset named RSVP-3M, featuring large-scale fine-grained instruction-following, is developed. In particular, diverse publicly available RS data are transformed and re-annotated into uniform conversation formats. Furthermore, part of the more high-quality caption data is generated from GPT-4V[18]. Those captions are uniquely tailored with the distinctive characteristics of each RS imagery, thereby enhancing the richness and diversity of data. Through the data conversion and re-annotation from existing datasets and GPT-4V, over 3M image-point-text and image-region-text pairings are constructed, covering a wide geographic distribution and multiple types of ground targets.
Extensive tests are conducted on multi-type RS datasets to evaluate the performance of EarthMarker which is demonstrated to be superior to state-of-the-art (SOTA) specialist models, MLLMs, and visual prompting models in various RS visual tasks at different granularity. Specifically, for the zero-shot scene classification task, EarthMarker shows a significant improvement compared with other existing MLLMs. Notably, for referring object classification, EarthMarker achieves a Semantic Similarity (SS) score of 98.37 using bounding boxes as visual prompts and 95.96 using point prompts on DIOR-RSVG dataset[19]. Furthermore, for image and region captioning tasks, EarthMarker also far exceeds other MLLMs and visual prompting models on the NWPU-Captions[20] dataset. In summary, the experimental results demonstrate that EarthMarker exhibits exceptional performance across a variety of multi-granularity RS image comprehension tasks and excellent zero-shot reasoning capability.
Our contributions can be summarized as follows.
The First RS Visual Prompting Dataset, RSVP. A large-scale RS regional instruction dataset named RSVP-3M, containing over 3M image-point-text and image-region-text pairings, is constructed. The construction of RSVP-3M facilitates fine-grained RS imagery interpretation, laying the foundation for the development of visual prompting in the RS domain.
The First RS Visual Prompting MLLM, EarthMarker. Leveraging our newly constructed RSVP, the visual prompting MLLM named EarthMarker is proposed. EarthMarker can interpret RS imagery in the multi-turn conversation at different granularity, including image, region, and point levels, significantly catering to the fine-grained interpretation needs for RS imagery.
The First RS Visual Prompt Learning Framework. A universal region and point-level visual prompting data annotation method is developed. Subsequently, a sharing visual encoding mechanism is proposed, which adapts visual prompts to match the dimensions of the input image, thereby both of them undergo uniform processing by the same visual encoder. This mechanism comprehensively enhances the interplay among visual prompts, holistic images, and text instructions. Furthermore, the cross-domain phased learning strategy is designed, and the disjoint parameters are optimized in a lightweight manner by leveraging the multi-domain data, endowing EarthMarker with spatial perception and visual prompting following capabilities.
Superior performance on multi-granularity RS Visual Tasks. Extensive experiments are conducted to demonstrate EarthMarker’s competitive performance in multi-granularity RS visual interpretation tasks, compared with the SOTA specialist models, MLLMs, and visual prompting models. The tasks evaluated include scene classification, referring object classification, captioning, and inter-relationship analyses. Therefore, EarthMarker successfully explores the adaptation of the visual prompt learning framework in the RS domain, improving the performance of MLLM and representing a significant step in fine-grained RS imagery interpretation.
II Related Work
II-A Multi-modal Large Language Models (MLLMs)
Recently, the advancement of large language models (LLMs) has significantly fueled the revolution and innovation in the natural language processing (NLP) field. The representative works including closed-source GPT series [21, 11] and open-source LLaMA series [22, 23] have achieved powerful generalizable language processing and reasoning ability. Inspired by LLM and by further injecting visual signals, MLLMs are developed for visual-language mutual comprehension and various visual tasks. For example, VisualGPT [24], BLIP [25] and Flamingo [26] show strong multi-modal reasoning potential after aligning LLMs with visual modality. Notably, LLAMA-Adapter V2 [27] and SPHINX [28] adopt zero-shot attention mechanism and linear projection layers tuning to mix LLM with visual signal. Those nature scene MLLMs laid the foundation for the extension to the remote sensing (RS) domain.
Some pioneer RS MLLMs have emerged, and related studies such as EarthGPT [5], Geochat [4], and SkyEyeGPT [29] have enabled MLLMs to interpret RS imagery. Among them, Geochat is the first MLLM targeting solving multiple tasks on optical RS images. Furthermore, EarthGPT has proposed a more universal MLLM that can deal with multi-source RS imagery and a wide range of RS visual tasks. There is no doubt that those models facilitate the development of MLLMs in the RS-specific domain. However, those models complete visual interpretation only through human-like language interactions, but cannot generate responses through visual prompts. Apparently, existing RS MLLMs mainly focus on image-level and visual grounding, but are incapable of referring comprehension. Therefore, this paper aims to enhance the MLLMs for referring fine-grained understanding of vision.
II-B Prompt Engineering
Prompt engineering is an emerging research direction in NLP [21]. Representation works contain AutoPrompt [7] and CoOp [8], which are designed to automate prompt template generation for language and vision-language models, instead of manual crafting. Additionally, Language prompting has been applied for developing open-vocabulary detection models such as DetPro [30] and Promptdet [31]. Compared with the extensively developed language prompting technique, visual prompting also needs more exploration. A major development is the Segment Anything (SAM) [10] model, which supports multiple segmentation prompts to enhance the zero-shot performance. Due to the lack of semantic labels in SAM, the Semantic-SAM [32] is proposed to realize multi-level semantics analysis and prediction. Notably, GPT4RoI [11] uses spatial boxes, and combines language and region-of-interest for input, enabling regional recognition. Colorful Prompting Tuning (CPT) [33] uses color-based markers to improve the performance of pre-trained vision-language models. The aforementioned models are trained on nature scene datasets. Note that Osprey [13] incorporates fine-grained mask regions into language instruction, achieving pixel-level visual understanding. Other visual prompting works including RegionBlip [12], Kosmos-2 [34], Shikra [35], and Ferret [14], also have shown promising results in region-based image understanding by leveraging visual prompting techniques. Additionally, the study entitled “Visual Prompting via Image Inpainting”[1] shows that various vision tasks can be accomplished well by giving desired task examples.
There are pioneering studies in the RS domain on region-level image understanding. For example, RSVG [16] can provide the referred object’s bounding box based on images and natural language expression. Moreover, EarthGPT [5] also has the visual grounding ability, and it is capable of providing captions for specific areas within images. Inspired by prompt learning, RSPrompter [17] designs an automated approach to generate appropriate prompts for SAM input, facilitating RS imagery segmentation. However, RSVG adopts language prompting but without visual prompting, whilst RSPrompter is only tailored to the segmentation task. Apparently, there is no unified visual prompting framework designed for the RS domain to further improve the performance of MLLMs. Those limitations hamper the development of more complex and fine-grained RS imagery understanding, therefore this paper focuses on filling this gap.
III Methodology
We first overview the overall model architecture in Section III-A. Subsequently, the three-phase continuous training strategy of the proposed EarthMarker is detailed in Section III-B.
III-A Model Architecture
One challenge in the RS domain is the absence of a visual prompts learning framework to endow MLLMs with fine-grained image understanding capabilities, blocking more complex reasoning. To address this challenge, EarthMarker is proposed, utilizing visual prompting for multi-granularity RS imagery comprehension. As illustrated in Fig. 2, EarthMarker contains four core components: a sharing visual encoding mechanism, a modality-align projection layer, a text tokenizer module, and a LLM decoder. These components work together to deal with multi-modal information, such as text instruction, images, and diverse visual prompting marks including bounding boxes and points, allowing LLM to generate accurate text responses. Each part is introduced as follows in detail.
In particular, the images and corresponding visual prompts share a visual encoding mechanism for feature sharing, enabling the visual encoders to better understand and associate the relationship between images and visual prompts. Specifically, the Mixture of Visual Experts (MoV) [36] is designed to encode the visual information. The MoV incorporates two visual encoders, DINOv2-ViT L/14 [37] and CLIP-ConvNeXt [38], which are pre-trained on distinct network architectures (ViT and CNN), thus offering complementary visual semantics. To refine the robust multi-scale visual features, the input images are downsampled to different resolutions denoted as and then fed into the MoV module to encode, respectively. Leveraging the strengths of various visual backbones, visual perception is enhanced and key details in images are refined. Subsequently, the encoded visual features are transformed to the same dimension and concatenated channel-wisely to obtain the integrated multi-scale feature maps represented as . This process can be formulated simply as
Notably, a key step to the encoder-sharing mechanism is the “Visual Prompt as Images”. Especially, the dimension () of the visual prompts is processed to the same dimension () with the images. Then, the transformed visual prompts also can be fed into MoV together with the images, the encoded visual prompts expressed as . Similarly, this process is written as
Subsequently, the modality alignment projection layer transforms the visual tokens into the language semantic space. Meanwhile, the text instructions are processed by the tokenizer module, which handles text tokenization and embedding, converting them into text embeddings . After obtaining the projected image tokens, visual prompts tokes, and text instructions embeddings, they are integrated into an entire multi-modal input sequence. LLM decoder takes the multi-modal inputs and generates the response sequence Y. Formulated as
We employ Llama 2, a transformer-based decoder-only LLM, as the LLM decoder.
III-B Cross-domain Phased Training
To realize the fundamental image-level understanding, spatial perception, and region/point-level RS data interpretation ability, the cross-domain phased training method is designed. The entire training process is divided into three phases including multi-domain image-text alignment, spatial perception tuning, and RS visual prompting tuning stage. Throughout the training, we keep lightweight training and avoid expensive full-parameters tuning. Furthermore, the disjoint parameters strategy is proposed, namely, the updated parameters of each stage are different. This strategy is conducive to the step-by-step solid understanding of images, and naturally solves the interference between image-text understanding, visual prompting comprehension, and fine-grained instruction-following.
Multi-domain Image-text Alignment. The first phase employs a multi-domain image-text alignment strategy. In this stage, both natural and RS domain image-level data are leveraged for pre-training to bring visual and text knowledge into alignment within a high-dimensional feature space. This strategy enables EarthMarker to deeply understand the holistic semantics of images. Specifically, we utilize the natural scene caption dataset COCO Caption [39], alongside the RS image caption and scene classification subset from the newly constructed RSVP. During this training phase, multi-scale visual features and language representations are integrated into the LLM to develop image-level comprehension capabilities. The MoV module is kept frozen throughout the training, so as to concentrate on refining robust visual features. Only the alignment projection layer, which acts as the visual-language connector, undergoes parameter updates to enhance the multimodal capabilities of the proposed EarthMarker and ensure seamless integration of visual and textual information.
Spatial Perception Tuning. In the previous step, EarthMarker achieved image-level comprehension capability. In this step, to acquire spatial perception and object-level comprehension, the nature scene publicly available datasets, RefCOCO [40] and RefCOCO+ [41] are transformed into the instruction-following format. Throughout the training, the attention layers of LLM are unfrozen for aligning spatial region features with language embeddings. Specifically, LLM’s key module self-attention head is composed of key , query , and the value , which are transformed by several linear layers. The -th implementation equations can be expressed as follows
where represents multi-modal input. The parameters , , , , , and are updated during the training. Then, the -th single attention scores of and are calculated as
where is the dimensionality of the keys. In addition, the other modules are kept frozen.
RS Visual Prompting Tuning. The last stage focuses on accurately following user instructions and achieving complex region-level and point-level visual reasoning tasks. The MoV, alignment projection, and LLM are fixed. The LoRA method is adopted for tuning. We load the weights trained in the previous phase and continue training EarthMarker on RSVP-3M region-text and point-text parings, which contain the fine-grained referring object classification and referring brief caption data. Specifically, several learnable low-rank adapter matrices are inserted into Transformer layers of LLM. The adapted multi-head attention is denoted as , the output of the -th adapted Transformer attention is formulated as
In conclusion, the cross-domain phased training endows EarthMarker with various granular (e.g., image-level, point-level, and region-level) multimodal instruction capabilities in the RS domain. In the first multi-domain image-text alignment stage, the LLM is efficiently converted into an MLLM, which is capable of image-level understanding. Subsequently, by utilizing the nature scene referring datasets, EarthMarker is equipped with the fundamental spatial perception of images. This is beneficial for subsequent developments of referring ability in the RS domain. Furthermore, by leveraging the RS visual prompting datasets RSVP-3M, EarthMarker is endowed with image understanding at both the region level and point level. Notably, different field’s datasets are adopted for training, and enhancing open-vocabulary reasoning ability. Notably, during the whole training, our updatable parameters are disjoint, preventing interference between understanding images at different granularity and the capability to follow visual prompts.
IV RS Visual Prompting Dataset Construction
In this section, a visual prompting dataset named RSVP-3M is presented. RSVP-3M is the first visual prompting instruction dataset in the RS field, designed to advance image-level and fine-grained point-level, region-level RS MLLMs. Specifically, RSVP-3M contains over 3 million multimodal dialogue data with visual prompting marks. Those multi-granularity visual prompting data are restructured and cleaned from existing publicly available RS datasets. Furthermore, the GPT-4V [18] is employed for automatic annotation to construct a high-quality complex visual reasoning dataset[15]. A detailed explanation of the construction of the RSVP-3M dataset is introduced as follows.
IV-A Data Conversion and Annotation from Public RS Datasets
A part data of the dataset RSVP-3M is constructed by restructuring and relabeling existing RS datasets. A range of visual task types is covered, containing image classification, instance segmentation, object detection, image caption, and region caption, see Tab. I. The image-level, region-level, and point-level data are derived from different RS datasets. Firstly, image-level visual prompting data is converted from image classification and captioning datasets. For the two type datasets, image-level visual instructions are used, with the bounding box serving as the visual prompt to obtain the image’s category or brief caption. Subsequently, the region-level data is based on object detection datasets. The ground truth bounding boxes are used as visual prompts to guide the model to identify the object-level or region-level categories accurately. Additionally, the point-level data is transformed from segmentation datasets. For instance segmentation, the representative points extracted from masks corresponding to instances are used as point-level visual prompts. For semantic segmentation, each image is divided into patches, and the points are randomly sampled within each patch as the visual prompts, with the category retrieved from the corresponding segmentation map.
Task | Raw Data | Samples |
Image Captioning | NWPU-Captions [20] | 169,981 |
RSITMD[42] | 24,387 | |
Sydney-Captions [43] | 2,837 | |
Region Captioning | DIOR-RSVG [19] | 31,491 |
Scene Classification | NWPU-RESISC45 [44] | 94,500 |
OPTIMAL 31 [45] | 3,720 | |
RSD46 [46] | 350,685 | |
WHU-RS19 [47] | 3,015 | |
Referring Object Classification (Box) | DOTA V2 [48] | 99,774 |
FAR1M [49] | 17,074 | |
NWPUVHR10 [50] | 1,888 | |
RSOD [46] | 1,465 | |
UCAS-AOD [51] | 3,203 | |
VisDrone [52] | 128,531 | |
MAR20 [53] | 2,096 | |
DOSR [54] | 1,127 | |
LEVIR [55] | 4,633 | |
HRSC2016 [56] | 2,171 | |
HRRSD [57] | 12,519 | |
Referring Object Classification (Point) | Vaihingen | 45,104 |
Potsdam | 504,139 | |
Hi-UCD [58] | 125,908 | |
NWPUVHR10 [50] | 956 | |
isAID [59] | 294,355 | |
Uavid [60] | 1,167,543 | |
SOTA [61] | 171,583 | |
FAST [61] | 263,056 | |
WHU [62] | 33,022 | |
Relationship Analyses | HRRSD [57] | 11,602 |
RSOD [46] | 934 | |
NWPUVHR10 [50] | 1,950 | |
LEVIR [55] | 3,956 | |
UUCAS-AOD [51] | 2,892 | |
DIOR-RSVG [19] | 15,274 | |
Grounded Captioning | DIOR-RSVG [19] | 15,274 |
GeoChat [4] | 35,264 | |
NWPUVHR10 [50] | 975 |
In RSVP-3M, each data item consists of visual prompts, user instructions, and images. The visual prompts in the user instructions or model answers are expressed as or . For the point-level data, for example, the user instruction which guides referring object classification is “Please identify the category of each marked point in the image”. The answer format is “”. In addition, for the region-level data, Take the scene of the airport as an example. The user instruction for airport region captioning is “Please provide the brief caption of each marked region in the image, and the corresponding answer format generated by the model is “ ”. The data structures of other visual tasks are similar to those explained above. Through the transformation and re-annotation based on public datasets, the visual prompting dataset RSVP-3M is effectively developed, featuring image-point-text and image-region-text pairings.
IV-B GPT4V-assisted Visual Prompting Data Generation
The aforementioned public datasets only provide simple classification information and brief captions, which are insufficient for intelligent interpretation of complex RS imagery. To mitigate the limitation and develop a more detailed and explicit RS visual prompting dataset, the language prompts for GPT-4V are carefully crafted for generating data featuring various complex visual reasoning. The complex fine-grained visual tasks involve detailed image captioning, inter-relationship analysis, and grounding captioning. We adopt the Set-of-Marks[63] (SoM) prompting, which can effectively unleash the extraordinary visual grounding ability of GPT-4V, to obtain comprehensive and unique characteristics from the RS imagery.
The data generated using GPT-4V not only compensates for the lack of information in brief captions but also provides detailed descriptions that reveal the spatial and semantic relationships between different regions in the image. For example, in aerial imagery, it is feasible to identify the general category of the image and provide a simple description. Additionally, detailed descriptions, such as the spatial layout of tennis courts, basketball courts, and playgrounds, the relationships among these areas, and the activities of people on the playground, can be conducted. The RSVP-3M dataset, supplemented by public datasets and data generated by GPT-4V, covers a wide range of fine-grained visual reasoning tasks, enhancing the richness and diversity of the data. The structure of this subset’s data is as follows.
V Experiments
In this section, we present extensive experiments to validate the superior performance of EarthMarker. In Section V-A, we introduce the implementation details. Subsequently, we conduct qualitative and quantitative analyses to provide a holistic view of EarthMarker’s performance from Sections V-B to V-E.
V-A Implementation Details
The proposed EarthMarker adopts the cross-domain phased training strategy, and the parameters updated vary at different stages. In general, we train an off-the-shelf 13B language model Llama 2 and the visual encoder MoV is kept frozen during the training. In the first multi-domain image-text alignment stage, only the alignment projection layer is updated. Then, in the spatial perception tuning phase, only the attention layers of LLM are unfrozen. Furthermore, the trainable LoRA metrics are introduced in the last RS visual prompting tuning stage. We utilized AdamW optimizer[64] with weight decay = 0 and betas = (0.9, 0.95), the learning rate is set to 2e-5, and the total training stages are conducted on 8 NVIDIA A100 GPUs. For model evaluation, we select diverse multi-granularity visual tasks to assess the performance of EarthMarker. Image-level tasks include scene classification, image captioning, and region-level tasks contain referring object classification and region captioning.
Method | Publication Year | BLEU1 | BLEU2 | BLEU3 | BLEU4 | METEOR | ROUGE-L | CIDEr(0-5) | SPICE |
Expert Model | |||||||||
CSMLF [69] | GRSL 2019 | 77.0 | 64.9 | 53.2 | 47.1 | 32.0 | 57.8 | 106.5 | 26.5 |
Qu, et.al [43] | CITS 2016 | 72.5 | 60.3 | 51.8 | 45.5 | 33.6 | 59.1 | 117.9 | 27.6 |
Attention(soft) [70] | TGRS 2017 | 73.1 | 60.9 | 52.5 | 46.2 | 33.9 | 59.9 | 113.6 | 28.5 |
Attention(hard) [70] | TGRS 2017 | 73.3 | 61.0 | 52.7 | 46.4 | 34.0 | 60.0 | 110.3 | 28.4 |
FC-Att [71] | TGRS 2019 | 73.6 | 61.5 | 53.2 | 46.9 | 33.8 | 60.0 | 123.1 | 28.3 |
SM-Att [71] | TGRS 2019 | 73.9 | 61.7 | 53.2 | 46.8 | 33.0 | 59.3 | 123.6 | 27.6 |
MLCA-Net [72] | TGRS 2022 | 74.5 | 62.4 | 54.1 | 47.8 | 33.7 | 60.1 | 126.4 | 28.5 |
Visual Prompting Model | |||||||||
EarthMarker(Ours) | 84.4 | 73.1 | 62.9 | 54.3 | 37.5 | 70.0 | 162.9 | 26.8 |
V-B Scene Classification
For scene classification tasks, we use the AID[73] and UCMerced[74] datasets for evaluation. AID is a large-scale aerial dataset collected from Google Earth, containing 30 categories. Following the setting of GeoChat, we use a 20 split of the AID dataset for testing. The UCMerced dataset consists of 21 categories for scene classification. Following the setting of GeoChat, the entire UCMerced dataset is adopted as a zero-shot test set.
We prompt the model with an image-level box to represent the entire image. The text instruction is “Please identify the object category of each marked region in the image.”. We calculate the zero-shot accuracy on the AID and UCMerced dataset. EarthMarker significantly outperforms other VLMs, with an accuracy of 86.52 on UCMerced and 77.97 on AID, as presented in Tab. II. In comparison, LLaVa-1.5 and Sphinx, due to the lack of RS domain knowledge, are inferior to the RS MLLM GeoChat and our EarthMarker. Compared to GeoChat, our EarthMarker achieved an accuracy improvement of 5.94 on AID and 2.09 on UCMerced.
Method | Publication Year | Format | BLEU-1 | BLEU-2 | BLEU-3 | BLEU-4 | METEOR | ROUGE | CIDER | SPICE |
MLLM | ||||||||||
Qwen-VL-Chat [68] | Arxiv 2023 | Coor | 20.66 | 9.55 | 4.90 | 1.96 | 8.35 | 20.93 | 29.18 | 9.55 |
GeoChat [65] | CVPR 2024 | Coor | 20.86 | 9.63 | 5.43 | 3.25 | 12.94 | 26.98 | 30.92 | 24.97 |
Sphinx [28] | Arxiv 2023 | Coor | 43.32 | 33.58 | 27.58 | 22.81 | 21.70 | 47.81 | 235.07 | 38.12 |
EarthGPT [5] | TGRS 2024 | Coor | 49.73 | 39.24 | 32.50 | 26.79 | 24.09 | 47.87 | 232.79 | 38.29 |
Visual Prompting Model | ||||||||||
Sphinx-V [15] | Arxiv 2024 | Box | 45.19 | 34.83 | 28.59 | 23.64 | 24.34 | 50.54 | 235.09 | 43.53 |
Vip-LLava-7b [75] | CVPR 2024 | Box | 21.03 | 11.26 | 6.37 | 3.25 | 9.59 | 23.50 | 30.95 | 12.44 |
Vip-LLava-13b [75] | CVPR 2024 | Box | 21.68 | 11.17 | 6.23 | 3.38 | 9.18 | 23.22 | 28.58 | 10.68 |
GLaMM [75] | CVPR 2024 | Box | 23.30 | 14.18 | 8.78 | 5.15 | 11.24 | 30.36 | 64.08 | 17.99 |
EarthMarker(Ours) | Point | 57.45 | 49.23 | 43.88 | 39.46 | 32.26 | 61.51 | 400.76 | 61.03 | |
EarthMarker(Ours) | Box | 57.14 | 48.60 | 43.06 | 38.59 | 31.97 | 60.46 | 379.25 | 59.87 |
Method | Publication Year | Format | SS | SIOU |
MLLM | ||||
GeoChat [68] | CVPR 2024 | Coor | 79.59 | 68.80 |
Sphinx [28] | Arxiv 2023 | Coor | 93.72 | 89.37 |
EarthGPT [5] | TGRS 2024 | Coor | 94.64 | 90.16 |
Visual Prompting Model | ||||
Sphinx-V [15] | Arxiv 2024 | Box | 89.07 | 81.62 |
Vip-LLava-7b [75] | CVPR 2024 | Box | 72.56 | 55.94 |
Vip-LLava-13b [75] | CVPR 2024 | Box | 74.51 | 60.53 |
EarthMarker(Ours) | Point | 95.96 | 93.49 | |
EarthMarker(Ours) | Box | 98.37 | 97.24 |
V-C Image Captioning
To evaluate the image captioning capabilities, we use the NWPU-Captions[20] dataset to assess and compare EarthMarker against other expert models in the supervised setting. Created by Northwestern Polytechnical University, The NWPU-Captions dataset includes 31,500 aerial images and 157,500 sentences for RS image description. Following the protocol of MLCA-Net, we employ BLEU1, BLEU2, BLEU3, BLEU4, METEOR, ROUGE-L, and CIDErD as evaluation metrics. In the evaluation, we use the sentence “Please provide a brief caption of each marked region in the image.” as text instruction and a full-image box as the visual prompt. As shown in Tab. III, compared to other expert models, EarthMarker demonstrate improvements in BLEU1, BLEU2, BLEU3, BLEU4, METEOR, and ROUGE-L by 7.4, 8.2, 8.8, 6.5, 3.5, and 9.9, respectively, and a 36.5 improvement in CIDErD.
V-D Referring Object Classification
The referring object classification task aims to identify the category within the referring region in the image. The metrics used to evaluate this task are two semantic relevance indicators—Semantic Similarity (SS) and Semantic Intersection over Union (S-IOU) to assess a model’s classification ability. Closed-set testing is adopted on the test sets of object-level dataset DIOR-RSVG [19]. The text instruction is “Please identify the category of the marked region in the image”, which along with the bounding boxes are fed into LLM to predict the category of regions. Due to the former MLLM (e.g., GeoChat, Sphinx, and EarthGPT) only accepting images and text as input, the region prompt for those MLLMs are coordinates information contained in the text instructions.
As the results shown in Tab. V, EarthMarker achieves 95.96 in SS and 93.49 in S-IoU using point-level visual prompts, and 98.37 in SS and 97.24 in S-IoU based on box-level visual prompts on the DIOR-RSVG dataset. Both sets of results significantly outperform the SOTA method. Furthermore, EarthMarker surpasses the previous SOTA model EarthGPT by 3.73 in SS and 7.08 in S-IoU on the DIOR-RSVG dataset, demonstrating its robust capability in fine-grained box-level classification. Additionally, in Fig. 3, there are considerable differences in predictions of regional object categories by EarthMarker and other MLLMs, as well as visual prompting models. It is evident that when faced with complex geographical scenes and blurry tiny objects, EarthMarker’s predictions are significantly accurate.
V-E Region Captioning
For the brief region captioning, the test set of DIOR-RSVG [19] is employed. Specifically, we adopt boxes as the visual prompt and a text prompt, such as “Please provide a brief caption of each marked region in the image,” to prompt EarthMarker to concisely describe the content of the specified region using a brief caption. Similar to the image captioning task, metrics like BLEU1, BLEU2, BLEU3, BLEU4, METEOR, ROUGE-L, and SPICE are used to evaluate EarthMarker and other MLLMs and visual promoting models in region target understanding.
As displayed in Tab. IV, on the DIOR-RSVG test set, compared to other MLLMs such as Qwen-VL-Chat, GeoChat, SPhinx, EarthGPT, and visual prompting models such as Sphinx-V, ViP-LLava, and GLAMM, EarthMarker shows improvements of 7.72, 9.99, 11.38, 12.67, blue12.49, 7.92, 10.97, 165.67, and 17.50 in BLEU1, BLEU2, BLEU3, BLEU4, METEOR, ROUGE-L, CIDER and SPICE, respectively. Furthermore, we visualize the instances of EarthMarker and other models on the region captioning task, as shown in Fig. 4. In complex RS scenarios with numerous targets and extensive geographic coverage, EarthMarker can accurately identify and describe various specified targets, which are challenging tasks for other models.
V-F Complex Visual Reasoning
In this part, we present the qualitative experimental result of EarthMarker to demonstrate its proficiency in completing complex RS tasks such as key target inter-relationship analyses. The text instruction for the relationship analyses task is “Please analyze the relationship between all marked regions in the image.” As shown in Fig. 5, when faced with an airport scenario, four visual models provided different responses.
Specifically, the response generated by GPT-4V does not specify the exact categories of each marked region. Additionally, GPT4V incorrectly describes Region 1 as “Region 1 supports all other regions indirectly by housing vehicles and equipment necessary for ground operations, including baggage handling, maintenance, and possibly emergency services”, whereas Region 1 actually contains fuel storage tanks. The Vip-LLava model incorrectly identifies all regions as containing airplanes. Note that Sphinx-V correctly identifies the types of objects in each region and analyzes the internal relationship of some regions. For example, the Sphinx-V answer “they are positioned in close proximity to each other, suggesting they are within the same area of the airport. The tanks are similar in appearance and likely serve the same function.” However, it fails to provide a comprehensive analysis of the relationships among all four regions, and incorrectly states that the airplane in Region 4 is in motion.
In contrast, our EarthMarker delivers an exemplary response. It firstly summarizes the relationship of all marked regions representing different elements of an airport environment. It then analyzes functionally similar areas in detail, stating that “ and are both storage tanks, likely used for fuel or other liquids, and are similar in shape and size, suggesting they are part of the same system or facility. and are both commercial airplanes, indicating the area is used for aircraft operations. The positioning of the planes and the tanks suggests a functional airport with operational infrastructure for aircraft and fuel services.” This response accurately reflects the diverse functionalities within the airport, demonstrating superior comprehension and analysis compared to the other visual prompting MLLMs.
VI Conclusion
In this paper, the fine-grained MLLM called EarthMarker, the first visual prompting model specifically designed for the RS domain, is proposed. Moreover, the RS visual prompting instruction dataset called RSVP is constructed for the first time, facilitating the development of fine-grained RS imagery comprehension. Furthermore, the visual prompts learning framework is developed. In particular, the shared visual encoding method is developed to uniformly refine multi-scale visual features and visual prompt content, which is beneficial for comprehensively understanding the interplay between visual prompts and the holistic image. Subsequently, the referring areas in the input are replaced by the proposed hybrid representation before being fed into the LLM to instruct the model to comprehend referring areas and toward specific predictions. Employing the RSVP-3M and the visual prompt learning framework, EarthMarker is equipped with multi-granularity visual understanding capability at the image, region, and point levels, making it able to simultaneously perform comprehensive and intelligent analysis in real-world scenarios. In the future, we plan to incorporate a broader range of visual modalities into EarthMarker, enhancing its multi-source imagery comprehension capabilities. In addition, we plan to support free-form shapes as visual marks to adjust the referring granularity flexibly.